write formula electric field due to point charge.
1
Expert's answer
2020-09-17T14:29:47-0400

Let the point charge be Q. Next, we may put a charge q to the point with distance r from Q. Therefore, the electric force will be

"F=k\cdot\dfrac{Qq} {r^2} ." Here k is a constant equal to "9\cdot10^9\mathrm{Nm^2/C^2}."

The modulus of the electric field is

"E=\dfrac{F} {q} =k\cdot \dfrac{Q} {r^2}"

The direction of the field is towards the vector from Q if it is positive and to Q if it is negative.

So we may rewrite it as

"\vec{E} =k\cdot\dfrac{Q} {r^3} \cdot\vec{r}."

In case of non-vacuum, we should add a factor "\varepsilon"

"\vec{E} =k\cdot\dfrac{Q} {\varepsilon r^3} \cdot\vec{r}."